OpenStudy (anonymous):

i need help with the following problem:2.1q-4.7-1.8q=0.7q-4.8. can anyone help me work through this?

OpenStudy (anonymous):

Yes Goal: You want to combine like terms. What is a like term? Like terms are individual numbers that have like quantities. For instance, 2.1 q is a like term with 0.7q You want to simplify it by turning 2.1q and 0.7q into one number of (Coefficient)q Also, -4.7 and -4.8 are like terms. So the process: First simplify each side Left: 2.1q - 1.8q = .3q -4.7 = -4.7 Left side simplified: .3q - 4.7 Right: .7q - 4.8 is already simplified. Next, write the equation with both sides simplified: .3q - 4.7 = .7q - 4.8 Simplify like terms, one term at a time to prevent errors from rushing. -4.7 = .4q - 4.8 .1 = .4q Now divide .1 by .4 q = .25 I apologize if I had a math error, but I showed you the process so you should be able to fix it if I did. Always check your work.